OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL!) Please help! : ) If h(x) = {(2,1),(1,-3),(5,0),(-2,3)} and f(x) = 3x^2+2, evaluate (h of f)(1)

OpenStudy (anonymous):

f(1)= 3(1^2)+2=3+2=5 now (h of f)(x)=h(f(x)) therefore (h of f)(1)=h(f(1))=h(5)=0

OpenStudy (anonymous):

@shoplovepc

OpenStudy (anonymous):

Thank you so much!!